16 One mole of an ideal gas is heated at a constant pressure of 101300 N
$m^{2}$, from 273.2 K to 373 K. Calculate the work involved R = 8.314
1
Expert's answer
2017-06-01T12:05:10-0400
A= ∫_(V_1)^(V_2)〖PdV=P∆V=ϑR∆T〗=1∙8.314∙(373-273.2)=829.7 (J)